A spinner is divided evenly into eight parts that are numbered 1 to 8. A four-sided die is numbered from 1 to 4. What is the probability of rolling the die and spinning the spinner to get a sum of nine?
Answer by stanbon(75887)

A spinner is divided evenly into eight parts that are numbered 1 to 8. A four-sided die is numbered from 1 to 4. What is the probability of rolling the die and spinning the spinner to get a sum of nine?
--------------------
(die/spinner pairs):1/8,2/7,3/6,4/5
=======
Total number of random pairs: 4*8 = 32
----------------
P(sum is 9) = 4/32 = 1/8
=========================
Cheers,
Stan H.